# Natural solution of SUSY $\mu$ problem from modulus stabilization in modular flavor model

### v1

@@ 5.2 Natural $\mu_{eff}$ with modulus stabilization at medium $\Im\tau$
Although the scalar potential for modulus in general diverges for $\tau\rightarrow i\infty$, the introduction of nilpotent Goldstino $X$ superfield or multiple moduli/matter fields may alter the shape of the scalar potential at intermediate values of $\tau$, making it possible the stabilization of modulus at $\tau_{0}$ with medium $\Im\tau_{0}$ (for example, $\Im\tau_{0}\sim 5-10$). Detailed discussions on such a possibility will be given in our subsequent studies. In this work, we just concentrate on the implications of modulus stabilization at medium $\Im\tau$ on the SUSY $\mu$-problem. Note that the stabilized VEV $\tau_{0}$ (with medium $\Im\tau_{0}$) can be regarded to lie close to the $i\infty$ fixed point.
To obtain the suppressed effective $\mu$-parameter, in addition to the suppression factor from powers of $q^{1/24}$ or $(2\Im\tau)^{-1}$, we would also like to choose the modular function that vanishes at $\Im\tau=i\infty$. For example, the modular functions satisfy
@@
as well as the vanishing singlet components from their direct products, such as the modular functions $Y_{\bf 1}^{(12)}(i\infty)$,$Y_{\bf 1^{\prime}}^{(16)}(i\infty)$.
The asymptotic behavior of the previous modular forms near $i\infty$ within the fundamental domain can be calculated from Petcov:2023vws (for $z_{0}$ with $\Im z_{0}$ being sufficient large)
$Y_{\bf 1^{\prime}}^{(4)}(z_{0})$ $=$ $-12p\epsilon+96p^{4}\epsilon^{4}~{},$
$Y_{\bf 1^{\prime}}^{(8)}(z_{0})$ $=$ $-6p\epsilon-1230p^{4}\epsilon^{4}~{},$
$Y_{\bf 1^{\prime\prime}}^{(8)}(z_{0})$ $=$ $144p^{2}\epsilon^{2}-2304p^{5}\epsilon^{5}~{},$ (73)
@@
Let us suppose that the modulus could be stabilized at the imaginary axis (within the fundamental domain) with $z_{0}=it_{0}$ for medium large $t_{0}$, when proper SUSY breaking contributions with non-vanishing $f_{0}$ are taken into account.
The effective $\mu$-term from the bilinear term for $H_{u}H_{d}$ in (65), which satisfies $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$ for $62k_{3}\in 24\mathbb{Z}+$ 6, is given by
$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta(it_{0})^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{u}+k_{d})/2}~{},$ (75)
$\approx$ $\alpha\Lambda_{SUSY}\eta(it_{0})^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{r}-k_{3}+3)/2}~{},$
$\approx$ $\alpha\Lambda_{SUSY}e^{\frac{k_{3}\pi t_{0}}{6}}(2t_{0})^{-\frac{k_{3}}{2}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{\frac{k_{r}+3}{2}}~{}.$
with proper normalization factors for $H_{u},H_{d}$ from the Kahler potential.
@@
$\left[e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}\right]^{\frac{k_{3}}{2}}\ll 1.$ (76)
As the factor
@@
$(r_{H_{u}},k_{H_{u}})\sim(~{}{\bf 1^{\prime\prime}},~{}14),~{}~{}(r_{H_{d}},k_{H_{d}})\sim(~{}{\bf 1},~{}15)~{},$ (78)
and the choice of modular forms $Y_{\bf 1^{\prime\prime}}^{(8)}$ (or possibly $Y_{\bf 1}^{(12)}$, $Y_{\bf 1^{\prime}}^{(16)}$ etc with proper $k_{H_{u}}$ and $k_{H_{d}}$) with $k_{3}=-21$ to generate suppressed $\mu_{eff}$ parameter. In this case, the effective $\mu_{eff}$ parameter is
@@
$=$ $4.46{\alpha}\Lambda_{SUSY}e^{-29\pi t_{0}/6}(2t_{0})^{16}~{},$
with the global normalization factor ${\bar{N}_{Y_{\bf 1^{\prime}}^{(8)}}}=32.2794$.
It can be calculated that the value of $\mu_{eff}\lesssim 10^{-20}M_{SUSY}$ can be satisfied for $t_{0}\gtrsim 5.58$. If we choose the modular form from direct product of $(Y_{\bf 1^{\prime}}^{(4)})^{k+2}$ with larger $k\in N$, smaller $t_{0}$ is needed to get electroweak scale $\mu_{eff}$ parameter. Tiny value of $\mu_{eff}$ is an joint effect of suppression from powers of $q$ (or $(2\Im\tau)^{-1}$) and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
@@ 6 Conclusions
It is known that the solutions of SUSY $\mu$ problem involve the forbidden of the bare $\mu$-term via some symmetry and the re-generation of the effective $\mu$-term after the symmetry breaking. We propose that, the symmetry that responsible for the solution of the SUSY $\mu$-problem is a modular flavor symmetry. The bare $\mu$-term is forbidden by modular symmetry after assigning non-vanishing modular weights and/or non-trivial singlet representations for $H_{u},H_{d}$ superfields. Then, the effective $\mu$-term will be re-generated after the modulus field is stabilized, with the corresponding modulus VEV breaks the modular symmetry completely or partially (at the fixed points).
Ordinary single modulus stabilization mechanism can be affected by SUSY breaking contributions. The goldstino, which is necessarily present as a consequence of UV SUSY breaking and mixes with gravitino, can be promoted into a constrained nilpotent superfield $X_{nl}$ and provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly. We discuss the stabilization mechanism of a single modulus field with the presence of SUSY breaking contributions, described in the non-linear SUSY realization scheme with nilpotent goldstino $X_{nl}$. The presence of $X_{nl}$ contributions can slightly change the original stabilized modulus VEV to a value located in the vicinity of the fixed points. Natural value of $\mu_{eff}$, which is much suppressed relative to the SUSY scale, can be the consequence of series expansion of typical modular forms in terms of the tiny deviation parameter (away from the fixed point $\omega$), which is similar to the mechanism to naturally generate mass hierarchies in the vicinity of the fixed points. For modulus VEV with medium $\Im\tau$, the much suppressed $\mu_{eff}$ is the joint consequence of suppression from powers of $q$ [or $(2\Im\tau)^{-1}$], with the exponent of the power determined by the assignments of the modular weights for $H_{u}$ and $H_{d}$, and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.

### v3

@@ 5.2 Natural $\mu_{eff}$ with modulus stabilization at medium $\Im\tau$
When the modulus field is stabilized near $i\infty$, similar to the case where modulus VEV lies in the vicinity of $\omega$, the exponentiation power $l$ of deviation parameter $\epsilon$ in the suppression factor for the $H_{u}H_{d}$ bilinear can also be proven Novichkov:2021evw to depend only on the $\mathbb{Z}^{T}_{3}$ charge of $H_{u},H_{d}$, rather than the explicit form of modular forms. Since the coefficient of $\epsilon^{l}$ cannot be simply determined and may involve additional suppression factors, we opt to directly calculate the suppression factor and the coefficient for the bilinear $H_{u}H_{d}$ using an explicit model construction.
[…]
To obtain the suppressed effective $\mu$-parameter, in addition to possible suppression factor from powers of $q^{1/24}$ (from $\eta(\tau)$) or $(2\Im\tau)^{-1}$ (from normalization factor), we would also like to choose the modular function that vanishes at $\Im\tau=i\infty$. For example, the modular functions satisfy
@@
as well as the vanishing singlet components from their direct products, such as the modular function $Y_{\bf 1}^{(12)}(i\infty)$.
The asymptotic behavior of the previous modular forms near $i\infty$ within the fundamental domain can be calculated from Petcov:2023vws (for $z_{0}$ with $\Im z_{0}$ sufficiently large)
$Y_{\bf 1^{\prime}}^{(4)}(z_{0})$ $=$ $-12p\epsilon+96p^{4}\epsilon^{4}+{\cal O}(\epsilon^{7})~{},$
$Y_{\bf 1^{\prime}}^{(8)}(z_{0})$ $=$ $-6p\epsilon-1230p^{4}\epsilon^{4}+{\cal O}(\epsilon^{7})~{},$
$Y_{\bf 1^{\prime\prime}}^{(8)}(z_{0})$ $=$ $144p^{2}\epsilon^{2}-2304p^{5}\epsilon^{5}+{\cal O}(\epsilon^{8})~{},$ (82)
@@
Let us suppose that the modulus field could be stabilized at the imaginary axis (within the fundamental domain) with $z_{0}=it_{0}$ for medium large $t_{0}$, when proper SUSY breaking contributions with non-vanishing $f_{0}$ are taken into account. Naively, if the desired suppression factor $10^{-20}$ to regenerate $\mu_{eff}$ comes solely from $\epsilon^{2}$ when $\Lambda_{SUSY}$ is naively taken to be the Planck scale, very large $t_{0}$ (here $t_{0}\gtrsim 10$) is needed. Such a large $t_{0}$ possibly causes difficulties for modulus stabilization realizations, as the scalar potential for modulus field generally diverges for $\tau\rightarrow i\infty$. Therefore, it is desirable to introduce additional suppression factors other than $\epsilon$ so that medium large $t_{0}$ can lead to desired suppression factors.
The effective $\mu$-term from the bilinear term for $H_{u}H_{d}$ in (74), which satisfies $k_{H_{u}}+k_{H_{d}}=k_{r}-k_{3}+3$, is given by
$\mu_{eff}$ $=$ $\alpha\Lambda_{SUSY}\eta(it_{0})^{-2k_{3}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{(k_{H_{u}}+k_{H_{d}})/2}~{},$ (84)
$\approx$ $\alpha\Lambda_{SUSY}e^{\frac{k_{3}\pi t_{0}}{6}}(2t_{0})^{-\frac{k_{3}}{2}}Y_{\bf r}^{(k_{r})}(it_{0})(2t_{0})^{\frac{k_{r}+3}{2}}~{},$
where proper normalization factors for $H_{u}$,$H_{d}$ from the Kahler potential are taken into account.
@@
$\left[e^{\frac{\pi t_{0}}{3}}(2t_{0})^{-1}\right]^{\frac{k_{3}}{2}}\ll 1,$ (85)
is tiny to act as an additional suppression factor other than $\epsilon$. As the factor
@@
$(r_{H_{u}},k_{H_{u}})\sim(~{}{\bf 1},~{}16),~{}~{}(r_{H_{d}},k_{H_{d}})\sim(~{}{\bf 1^{\prime}},~{}16)~{},$ (87)
and the choice of modular form $Y_{\bf 1^{\prime\prime}}^{(8)}$ (or, for example, $Y_{\bf 1}^{(12)}$, $Y_{\bf 1^{\prime}}^{(16)}$ etc with proper $k_{H_{u}}$ and $k_{H_{d}}$) with $k_{3}=-21$ to generate a suppressed $\mu_{eff}$ parameter. In this case, the effective $\mu_{eff}$ parameter is
@@
$=$ $7.05{\alpha}\Lambda_{SUSY}e^{-29\pi t_{0}/6}(2t_{0})^{16}~{},$
where ${\bar{N}_{Y_{\bf 1^{\prime\prime}}^{(8)}}}=20.4152$ is the global normalization factor for the modular form $Y_{\bf 1^{\prime\prime}}^{(8)}$.
Calculations show that the value $\mu_{eff}\lesssim 10^{-20}M_{SUSY}$ can be achieved for $t_{0}\gtrsim 5.58$, if $\Lambda_{SUSY}$ is naively taken to be the Planck scale. For $\Lambda_{SUSY}$ typically of the order $\sqrt{F}\sim\sqrt{m_{3/2}M_{Pl}}\sim 10^{10}$ GeV, it is sufficient to choose $t_{0}\gtrsim 3.2$ to satisfy $\mu_{eff}\lesssim 10^{-8}M_{SUSY}$. The tiny value of $\mu_{eff}$ results from the combined effect of suppression from powers of $q$ (or $(2\Im\tau)^{-1}$) and the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$.
It can also be verified that the suppression power of $\epsilon=\exp[-\frac{2\pi}{3}\Im z_{0}]$ for $H_{u}(~{}{\bf 1},~{}16)$ and $H_{d}(~{}{\bf 1^{\prime}},~{}16)$ bilinear is given by the power factor $l$ of $(\rho_{i}^{c}\rho_{j})^{*}=\zeta^{l}$ for $\zeta=\exp(2\pi i/3)$, which can be read out from Novichkov:2021evw to be $(1\cdot\zeta)^{*}=\zeta^{2}$ and predict $l=2$. This power factor for $\epsilon$ is consistent with the result from our explicit construction.
@@ 6 Conclusions
In this study, we propose a novel solution to the SUSY $\mu$-problem within the framework of modular flavor symmetry. The explicit $\mu$-term is prohibited by modular symmetry after assigning non-vanishing modular weights and/or non-trivial singlet representations to the $H_{u},H_{d}$ superfields. The effective $\mu$-term is then regenerated following the stabilization of the modulus field, with the corresponding modulus VEV breaks the modular symmetry partially (at the fixed points) or completely (in the vicinity of the fixed points). Since naturalness requires $\mu_{eff}$ to lie at the electroweak scale or soft SUSY breaking scale, a large suppression factor leading to $\mu_{eff}\ll\Lambda_{SUSY}$ is required when the modulus field is stabilized. Such a large suppression factor can naturally arise from a modulus VEV in the vicinity of certain fixed points.
Ordinary single modulus stabilization mechanism can be affected by SUSY breaking contributions. The Goldstino, which is necessarily present as a consequence of UV SUSY breaking and mixes with gravitino, can be promoted into the constrained nilpotent superfield $\textbf{X}_{nl}$ and provides a way for an arbitrary non-supersymmetric low-energy theory to realize SUSY non-linearly. We discuss the stabilization mechanism of a single modulus field in the presence of SUSY breaking contributions, described by the non-linear SUSY realization scheme with a nilpotent Goldstino $\textbf{X}_{nl}$. The presence of $\textbf{X}_{nl}$ contributions can slightly alter the original stabilized modulus VEV to a value near the fixed points.
Natural value of $\mu_{eff}$, which is significantly suppressed relative to the SUSY scale, can be achieved from the expansion of typical modular forms using a tiny deviation parameter near the fixed points. This is similar to the mechanism that can naturally generate mass hierarchies for quarks and leptons in the vicinity of certain fixed points. The exponentiation power of the deviation parameter in the suppression factor, in fact, depends solely on the $\mathbb{Z}_{3}^{ST}$ charge (for fixed point $\omega$), or $\mathbb{Z}^{T}_{3}$ charge (for fixed point $i\infty$) of $H_{u}$, $H_{d}$, rather than the explicit form of modular forms.
For a modulus VEV with an intermediate value of $\Im\tau$, the significantly suppressed $\mu_{eff}$ results from the combined effects of suppression by powers of $q^{1/24}$ [or $(2\Im\tau)^{-1}$], where the exponentiation power determined by the modular weights of $H_{u}$ and $H_{d}$, along with the asymptotic suppression behavior of typical modular forms away from the fixed point $i\infty$, taking the form of appropriate power of the tiny deviation parameter.
The modulus stabilization mechanism with nilpotent $\textbf{X}_{nl}$ superfield can be fairly generic and ubiquitous. The regeneration of small $\mu_{eff}$ term following modulus stabilization, which just needs a modulus VEV near some fixed points, is independent on the concrete modulus stabilization mechanism.